A system of lines is given as y=mix+ci,where mi can take any value out of 0,1,-1 and when mi is positive ,then ci can be 1 or -1 when mi equal 0 or 1 and when mi equals to -1,ci can take 0 or 2. Then the area enclosed by all these straight lines is

Grade:

2 Answers

Saurabh Koranglekar
4 years ago
Dear student

Please check

when mi equal 0 or 1 and when mi equals

In the question.

Kindly repost with correction

Regards
Sia
3 years ago
According to question,we get 6 equations that are y=x+1,y=x-1,y=0,y=1,y=-x,y=2-x
If we draw all these lines in a graph we get hexagon with points (-1/2,1/2),(0,1),(1,1),(3/2,1/2),(0,0),(1,0) and if we find are we get 1/4+1+1/4=3/2